OpenStudy (kpop4life123):

I WILL FAN AND MEDAL!!!!!!!!!!!!!!!!!!!!!!!! There are (10^8)^2 ⋅ 10^0 candies in a store. What is the total number of candies in the store?

OpenStudy (campbell_st):

what do you know about anything to the power of zero

OpenStudy (kpop4life123):

im bad at math i dont acaully know

OpenStudy (campbell_st):

and for the 1st part you need to know about power of a power \[(x^a)^b = x^{a \times b}\] multiply the powers... so you need to use this index law as well

OpenStudy (campbell_st):

ok... so the index law for power of zero is \[a^0 = 1\] or when a base is raised to a power of zero, the value is 1